ishanii:

Alan weighs 3 pounds more than twice what Dylan weighs. Write an expression to represent Alan’s weight in terms of Dylan’s weight. 15px

Shadow:

So basically here you're just translating the sentence into the math that it's talking about. Let's start by labelling Alan's weight as variable y and Dylan's weight as variable x. Alan weighs translates to basically \[y =\]

Shadow:

Do you have an idea of where we would go from there?

ishanii:

its hard but are we trying to find X

Shadow:

Well in this case the topic of the sentence is Alan's weight, which we labeled y.

Shadow:

But we aren't actually trying to find y or x, we're just trying to translate what the sentence is saying, into math. So when it says "Alan weighs three pounds more" what do you think that would mean in math?

Shadow:

Here's an example. The cat weighs four less pounds than the dog. Let y be the cat's weight and x be the dog's weight. y = x - 4

ishanii:

you would need to make an equation

Shadow:

Yes, we're in the process of creating that equation. So how would you mathematically represent: "Alan weighs three pounds more"

Shadow:

If you are confused, try reading my example

ishanii:

so dylan weighs 6 pounds

ishanii:

so i mixed them up

ishanii:

alan weighs 6 pounds

Shadow:

Technically we're not trying to find Dylan's weight, we're just trying to create the equation that represents their weights.

ishanii:

oh i got it

ishanii:

2x+3=Alan's Weight

ishanii:

cause 2 times aka 2x times 3 is 6

Shadow:

Yeah. If X is dylan's weight then 'two times Dylan's weight' is 2x. Then three pounds more than 'two times Dylan's weight' is 2x + 3

Shadow:

I see how you're getting 6 but that actually isn't necessary nor correct. You can't compute 2x + 3 without knowing what x is.

ishanii:

oh i see

Shadow:

For example if x were 1, you could do 2(1) + 3 = 2 + 3 = 5

Shadow:

If x were 2 you could do 2(2) + 3 = 4 + 3 = 7

ishanii:

oh

ishanii:

i thought it was like an exponent

Shadow:

Yeah so the final equation would look like\[y = 2x + 3\]

Shadow:

Where, y = Alan's weight x = Dylan's weight

ishanii:

i didn't understand